OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9

OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 3x - 6y = 12 -2x + 3y = 6 -14 24 14 -24

OpenStudy (tyteen4a03):

@srijit Please don't give out answers. It doesn't help the asker in answering future questions like this. The correct way to do this are by either substituting or eliminating. I will show you how to substitute here as there's a bare y in there. 4x + y = 57 Subtract 4x from both sides and get y = 57 - 4x. Now substitute this fact into 7x - 2y = 21: 7x - 2(57 - 4x) = 21. Solve for x and there goes your answer. If other questions ask for y too, just substitute x into either equation to find y. The quick way: Plug in the values, and if any of them match the equation (i.e something = something) then it's the correct answer. Now solve the second question, good luck :)

OpenStudy (anonymous):

I kno how to do the add and subtract both sides an all dat i jus dont get how to put both in one equation

OpenStudy (tyteen4a03):

(Since the second question is faster by using elimination, here we go:) We know that: 3x - 6y = 12 -2x + 3y = 6 To eliminate, one of the variables must be of the same coefficient. Here I'll pick 3y as 2*3 = 6. Now multiply the whole equation -2x + 3y = 6 by the factor of 2 to get -4x + 6y = 12. Now the system looks like this: 3x - 6y = 12 -4x + 6y = 12 Now, add the two equation together to get -x = 24 (i.e x = -24). This method is quicker when you see 2 variables with common factors, but sometimes substitution is quicker if there's a bare variable with no coefficient, or if the coefficient of the variable is the same.

OpenStudy (anonymous):

@tyteen4a03 yea sure! :)

OpenStudy (tyteen4a03):

@KCB Suppose we have these equations: 1. x + 2y = 8 2. x + 8y = 64 Basically you want to move the terms around to get the desired one-variable equation: x/y = something. In this case, we'll use x as the subject, and we'll pick 1 as an example. (2 works too) x + 2y = 8. Simply subtract 2y from both sides to get x = 8 - 2y. Here is your intermediate equation. Now use this known fact and substitute to the other equation (so that the 2nd equation now becomes (8 - 2y) + 8y = 64 and find y. Then use the value of y to find x.

OpenStudy (anonymous):

ohh thank you so much i understand now :)